1)  The price has just increased from $1.50 to $3. Calculate the price elasticity of demand at $1.50. Round your answer to 2 decimal places if necessary.

 

 


 

2)   At a price of $1.50 is price elasticity of demand elastic, unit elastic or inelastic? Explain your answer.

 

 

3)   Calculate the total expenditure on lattes at price $1.50 per latte.

 

 

 

 


1
Expert's answer
2021-10-26T09:34:37-0400

1) "e_p= \frac{dQ}{dP}"



"e_p = \frac{40- 10}{(40 + 10)/2} * 100\%/\frac{1.50 - 3.00}{(1.50 + 3.00)/2} * 100\%"


"e_p = \frac{30}{25} * 100\%/\frac{-1.50}{2.25} * 100\% = \frac{1.2}{0.66} =1.818"



2) The price elasticity of demand is elastic since the percentage change in quantity is greater than the percentage change in price and the computed elastic is greater than "1"




3) Total expenditure equals the total area under the curve and below the price of $ 1.50


"= 1.50 * 40 + \frac{1}{2} * 30*1.50 = 82.5"